Given f(x)=x²+2x+1, find f(x+h) and simplify.
Answer by nyc_function(2741)

Let x = (x + h)

f(x + h) = (x + h)^2 + 2(x + h) + 1

f(x + h) = x^2 + 2xh + h^2 + 2x + 2h + 1

All you had to do is replace every x in the function with the value (x + h) and simplify.